# Strongly Interacting Dark Matter admixed Neutron Stars

### v1

@@ 1 Introduction
Due to the gravitational interactions, dark matter will necessarily accumulate in neutron stars. Thus, neutron stars make an excellent laboratory for testing such gravitational signals as their properties are determined by gravitational interactions Akmal:1998cf ; Lattimer:2000nx ; Ozel:2016oaf , and can be tested using direct observations and gravitational wave astronomy. However, it is a-priori unclear whether the quantity of accumulated dark matter is sufficient to produce observable effects. In particular, depending on the age, history, and location of a neutron star, the accumulated dark matter amount can vary wildly from anything upward from $10^{-3}$of the neutron star mass, but likely will exceed a percent only in the most extreme circumstances Guver:2012ba ; Biesdorf:2024dor ; Hajkarim:2024ecp ; Kunkel:2024otq ; Wystub:2021qrn ; Pitz:2024xvh ; Diedrichs:2023trk ; Dengler:2021qcq ; Tolos:2015qra ; Mukhopadhyay:2015xhs ; Rezaei:2016zje ; Panotopoulos:2017idn ; Nelson:2018xtr ; Ellis:2018bkr ; Gresham:2018rqo ; Ivanytskyi:2019wxd ; Karkevandi:2021ygv ; Sen:2021wev ; Guha:2021njn ; Goldman:2013qla ; Xiang:2013xwa ; Li:2012ii ; Sandin:2008db ; Leung:2011zz ; Leung:2012vea ; Barbat:2024yvi ; Hippert:2021fch .
@@
Fortunately, we do have a very good overview of global neutron star properties from observations. Among the direct observations, the Neutron star Interior Composition ExploreR (NICER) mission is dedicated to direct observation of neutron star properties via observations in the soft (0.2–12 keV) X-ray band Gendreau2024 . The mission concentrates on measurements of neutron star mass and radius, and will provide in its runtime a wide range of characterized neutron stars. Thus, a stochastically large enough sample is available to also detect variations of neutron star properties due to varying amounts of dark matter admixtures. In contrast to a single-component neutron star of ordinary matter, which yields a unique mass-radius relation, differing amounts of dark matter will fuzzy out such a curve. Thus, first-principle knowledge of this fuzzying, which we will present here, will allow to detect dark matter accumulation.
@@
Moreover, neutron stars with a large amount of dark matter in the interior, also known as mixed neutron stars, will provide potentially spectacular observations. Mixed neutron star binaries are thus a further testbed for dark matter properties, as they are capable of generating exotic signals at gravitational waves experiments. In particular, the so-called tidal deformability, indicating the deformation of colliding neutron stars, has been constrained by LIGO observations of a single binary neutron star merger GW170817 LIGOScientific:2017vwq ; LIGOScientific:2017ync . Along with this LIGO has also measured the “chirp mass” for the same system. These observations have been combined with simultaneous observations from other multi-messenger telescopes, leading to an even more comprehensive analysis LIGOScientific:2017ync .
@@
A primary ingredient in understanding such observable effects is the equation of state, which determines the evolution of thermodynamic properties, for both the ordinary component and a potential dark matter component inside the neutron star. Owing to the sign problem Gattringer:2010zz , lattice QCD, – the methodological mainstay to determine most of QCD physics – is unable to provide a fully controlled equation of state of ordinary matter inside neutron stars. Limited knowledge can be derived from heavy-ion experiments, models, and other methods Friman:2011zz . However, this has so far been insufficient to provide a complete understanding. In fact, observed neutron star properties are used to constrain the QCD equation of state Akmal:1998cf ; Lattimer:2000nx ; Ozel:2016oaf ; LIGOScientific:2017ync , under the assumption that neutron stars are primarily composed of ordinary matter. As a consequence, we will employ here three different candidates for the equation of state of ordinary matter, to represent this uncertainty. The particular choices will be discussed in section 2.
@@
Therefore, while neutron stars may be efficient dark matter accumulators, and consequently dark matter discovery avenues, insufficient knowledge on the accumulated amount coupled with limited control over neutron star ordinary physics present a challenging situation. Progress with this conundrum would therefore be made if at least the dark matter equation of state would be controlled from first principles, reducing the uncertainties on the evolution of one of the two neutron star components. This is possible for WIMPs (and similar) cases, and several studies have been performed in this regard Quddus:2019ghy ; Quddus:2020lbt .
@@
Therefore any study of mixed neutron stars involving a new QCD-like admixture will benefit from a dark QCD-like sector with fermionic dark hadrons, where the equation of state is accessible in first-principle calculations. Fortunately, such a possibility exists, e. g. , a hidden QCD-like sector based on the gauge group G2 Holland:2003jy ; Maas:2012wr ; Wellegehausen:2013cya . It has already been shown that this theory can support pure compact stellar objects Hajizadeh:2017jsw , which could be interpreted as pure dark neutron stars. Already the one-flavor theory provides a conserved quantum number carried by the fermionic dark baryons Maas:2012wr , and thereby offers a dark matter candidate, alongside a rich spectrum of further states Wellegehausen:2013cya . The two-flavor case furthermore offers dark pions as additional stable states with a Wess-Zumino-Witten-like $3\to 2$ process, relevant to the dark matter relic density generation mechanism Hochberg:2014dra ; Hochberg:2014kqa ; Kulkarni:2022bvh . There is also an additional $3\to 2$ process which converts dark pions into dark fermionic baryons, and allows to build Fermi pressure, from accrued bosonic dark matter, required for neutron star stabilization Hajizadeh:2017jsw . G2-QCD therefore provides an ideal testbed for the previously mentioned challenge.
@@
For ordinary matter, we use a selection of the currently favored equations of state Kurkela:2014vha ; Dengler:2021qcq ; Pitz:2024xvh covering a large range of phenomenology. The results will therefore only be quantitatively relevant if either we keep the amount of dark matter limited such that, within errors, neutron star properties remain the same. Or we stipulate that most neutron stars do not contain large amounts of dark matter, but some do, giving a distinct signal differing from ordinary neutron stars when observed. Alternatively, one can interpret our injection of G2-QCD equation of state seriously and demand modification of the ordinary matter equation of state to compensate for large modifications in the neutron star properties. However, this is beyond the scope of the present exploratory work, especially in view of the restricted set of available equations of state for the G2-QCD case. We will also assume throughout that the interactions between dark matter and ordinary matter, even at large densities of both, are negligible. This may be justified given that in most strongly-interacting dark matter scenarios the portal couplings to the SM are highly suppressed Hochberg:2015vrg ; Berlin:2018tvf ; Pomper:2024otb ; Bernreuther:2023kcg .
For the sake of completeness, we discuss briefly our choices of the equations of state in section 2, together with a brief overview of the salient features of G2-QCD. Our approach to building a two-component neutron star is standard Hajkarim:2024ecp , and briefly rehearsed in section 3. In section 4 we provide results for a range of dark matter masses and dark matter content. We provide the mass-radius relations, together with maximal and minimal values for both. We also provide the tidal deformability, which is the primary input to the gravitational wave signal for the merging of two neutron stars. Finally, taking our inputs at face value, we interpret our result in terms of limits on dark matter properties or dark matter contents in neutron stars. This part should be considered rather a blueprint than a hard statement, given our limited inputs, but can be improved arbitrarily by improving these inputs. Additional results are available in appendices 6.1 and 6.2.
@@ 2 Equations of state
The equation of state is the key ingredient for the Tolman-Oppenheimer-Volkoff (TOV) equation Oppenheimer:1939ne . To describe a neutron star consisting of two kinds of fluid we need the equations of state describing the two fluids.
@@ 2.1 Ordinary matter
We want to cover a large range of possible equations of state for describing the ordinary matter. To this end, we work with the equations of state from Kurkela:2014vha . This work uses piecewise polytropes, to interpolate between a low density regime described by nuclear chiral perturbation theory (NChPT) Hebeler_2013 and a high density regime described by perturbative QCD Kurkela:2009gj . Equations of state that do not meet the observational constraints from NICER data on the mass, radius and the tidal deformability are omitted. For this study, we use 3 limiting equations (EoS I, II and III) of state to get a complete picture Kurkela:2014vha . EoS I is the softest of the three barely reaching the two solar mass limits. EoS II yields a maximal maximum mass. EoS III is the stiffest and while it might be too stiff to describe an ordinary neutron star, we still include it in our analysis as the addition of dark matter to a neutron star usually results in smaller masses and radii. Therefore it might be able to meet observational constraints when considering mixed neutron stars. For the same reason we will employ central pressures larger than the ones stated in Kurkela:2014vha that result in the maximal mass. We will see that the addition of dark matter allows for stable solutions at larger central pressures.
@@ 2.2 G2-QCD
Details on the equation of state, its determination and properties, can be found in  Wellegehausen:2013cya ; Maas:2012wr ; Hajizadeh:2017jsw . Here, we concentrate on the equation of state at zero temperature, which has been discussed in Wellegehausen:2013cya ; Hajizadeh:2017jsw . Of course, the lattice data is only available at discrete values of the dark baryonic chemical potential. The interpolation between data points is, as for the ordinary matter equation of state, performed using piecewise polytropes to ensure thermodynamic consistency. Following Hajizadeh:2017jsw , the equation of state is extrapolated to small densities by a free Fermi fluid which follows $n=c(\mu-m_{C})^{3/2}$ where $m_{C}$ is the mass of the dark matter candidate. At high densities we extrapolate the data by a Fermi-Dirac distribution $n(\mu)=n_{s}/(\exp(a-b\mu)+1)$ fitted to the last few data points Wellegehausen:2013cya ; Hajizadeh:2017jsw . The upper bound for the equation of state is given by the point where the speed of sound reaches c. Herein $n_{s}$ is the lattice saturation density, i. e. the situation where Pauli blocking obstructs larger densities. This is a pure lattice artifact but the densities required in the following, as has already been observed previously Hajizadeh:2017jsw , are small enough for this to be not a relevant issue.
The equation of state has been obtained for two different values of the dark quark mass Wellegehausen:2013cya . The masses are chosen such that the lightest dark hadronic state – the dark pion, which is also unstable due to potential decays to the standard model via messengers – has a mass of 25% and 33% of the dark matter candidate Wellegehausen:2013cya .
@@ 3 Building a two-component neutron star
Figure 1: The equations of state from Kurkela:2014vha and Wellegehausen:2013cya used in this work. The dark matter candidate mass is chosen to be $m_{C}$ = 1 GeV for this plot for comparison. The end of the line indicate the maximum value for the pressure for the respective equation of state used in this work. The black line indicates $\varepsilon=p$.
@@ 3.1 Two-fluid TOV
The coupled two-fluid TOV equations can be solved using the equations of state of the two fluids. Iterating over the potentially nonidentical central pressures, one obtains the neutron star properties for each combination of central pressures ($p_{0,O}$ and $p_{0,D}$) by integrating the equations until the pressure vanishes. The radius at which the ordinary/dark matter pressure drops to 0 is labeled $R{O}$/$R{D}$. $M_{O}$/$M_{D}$ are the integrated masses of the two fluids respectively and $M_{tot}=M_{O}+M_{D}$. The solutions can result in a neutron star with a dark core ($r_{O}>r_{D}$) or a dark halo ($r_{O}<r_{D}$), which potentially also affects the gravitational wave signal of an inspiral of two neutron stars Ivanytskyi:2019wxd .
Dimensional analysis show that the mass and the radius scale like $1/m_{N}^{2}$. A hypothetical neutron star consisting of a 500 MeV neutron would therefore have 4 times the mass and the radius of a realistic neutron star. To some extent, this scaling argument is still present in the two fluid case. A lighter dark matter candidate mass will result in a stronger impact of dark matter on the star. In the limiting cases where the central pressure of one of either the dark or ordinary matter is negligible ($p_{0,D}\gg p_{0,O}$ or $p_{0,D}\ll p_{0,O}$) this scaling argument is perfectly recovered as very heavy dark matter candidates will have no impact on the observables while a light dark matter candidate will result in a compact object dominated by dark matter that can eventually not be interpreted as a neutron star anymore.

### v2

@@ 1 Introduction
Particle dark matter will necessarily accumulate in neutron stars making them an excellent laboratory to study dark matter gravitationally on small scales Akmal:1998cf ; Lattimer:2000nx ; Ozel:2016oaf . The amount of dark matter may vary depending on the age, history, and location of the neutron star, but most estimates conclude that that it is very unlikely to exceed 1% of the total neutron star mass Guver:2012ba ; Biesdorf:2024dor ; Hajkarim:2024ecp ; Kunkel:2024otq ; Wystub:2021qrn ; Pitz:2024xvh ; Diedrichs:2023trk ; Dengler:2021qcq ; Tolos:2015qra ; Mukhopadhyay:2015xhs ; Rezaei:2016zje ; Panotopoulos:2017idn ; Nelson:2018xtr ; Ellis:2018bkr ; Gresham:2018rqo ; Ivanytskyi:2019wxd ; Karkevandi:2021ygv ; Sen:2021wev ; Guha:2021njn ; Goldman:2013qla ; Xiang:2013xwa ; Li:2012ii ; Sandin:2008db ; Leung:2011zz ; Leung:2012vea ; Barbat:2024yvi ; Hippert:2021fch ; Issifu:2024htq ; Karkevandi:2024vov ; Shakeri:2022dwg ; Ema:2024wqr . Nevertheless, a careful analysis of potential effects of dark matter on gravitational and electromagnetic signals from neutron stars is necessary to assess if, and if yes what, impact dark matter could have on its properties. This will allow to understand whether neutron stars can be used to learn about dark matter. It also will clarify whether dark matter could, potentially at least, create effects relevant to understanding the ordinary matter aspects of neutron stars. Both aspects require a good data basis on neutron stars.
@@
On the observational side, fortunately, we do have a very good overview of global neutron star properties from observations. Among the direct observations, the Neutron star Interior Composition ExploreR (NICER) mission is dedicated to direct observation of neutron star properties via observations in the soft (0.2–12 keV) X-ray band Gendreau2024 . The mission concentrates on measurements of neutron star mass and radius, and will provide in its runtime a wide range of characterized neutron stars. Thus, a stochastically large enough sample is available to also detect variations of neutron star properties due to varying amounts of dark matter admixtures. In contrast to a single-component neutron star of ordinary matter, which yields a unique mass-radius relation, differing amounts of dark matter will fuzzy out such a curve. Thus, first-principle knowledge of this fuzzying, which we will present here, will allow to detect dark matter accumulation.
@@
Moreover, neutron stars with a large amount of dark matter in the interior, also known as mixed neutron stars, might provide smoking gun signals like supplementary peaks in the gravitational wave power spectral density Giangrandi:2022wht ; Ellis:2017jgp . Mixed neutron star binaries are thus a further testbed for dark matter properties, as they are capable of generating exotic signals for gravitational waves experiments. In particular, the so-called tidal deformability, indicating the deformation of colliding neutron stars, has been constrained by LIGO observations of a single binary neutron star merger GW170817 LIGOScientific:2017vwq ; LIGOScientific:2017ync . Along with this LIGO has also measured the “chirp mass” for the same system. These observations have been combined with simultaneous observations from other multi-messenger telescopes, leading to an even more comprehensive analysis LIGOScientific:2017ync .
@@
On the theoretical side, a primary ingredient in understanding such observable effects is the equation of state, which determines the evolution of thermodynamic properties, for both the ordinary component and a potential dark matter component inside the neutron star. Owing to the sign problem Gattringer:2010zz , lattice QCD, – the methodological mainstay to determine much of QCD physics – is unable to provide a fully controlled equation of state of ordinary matter inside neutron stars. Limited knowledge can be derived from heavy-ion experiments, models, and other methods Friman:2011zz . However, this has so far been insufficient to provide a complete understanding. In fact, observed neutron star properties are used to constrain the QCD equation of state Akmal:1998cf ; Lattimer:2000nx ; Ozel:2016oaf ; LIGOScientific:2017ync , under the assumption that neutron stars are primarily composed of ordinary matter. As a consequence, we will employ here three different candidates for the equation of state of ordinary matter, to represent this uncertainty. The particular choices will be discussed in section 2. However, they are made to reduce model dependence as much as possible in favor of data-driven input.
@@
Therefore, while neutron stars may be efficient dark matter accumulators, and consequently dark matter discovery avenues, insufficient knowledge on the accumulated amount coupled with no full control over neutron star ordinary physics present a challenging situation for the subject of dark matter admixed neutron stars. Progress with this conundrum would therefore be made if at least the dark matter equation of state would be controlled from first principles, reducing the uncertainties on the evolution of one of the two neutron star components. This is possible for WIMPs (and similar) cases, and several studies have been performed in this regard Quddus:2019ghy ; Quddus:2020lbt .
@@
Therefore any study of mixed neutron stars involving a new QCD-like admixture will benefit from a dark QCD-like sector with fermionic dark hadrons, where the equation of state is accessible in first-principle calculations. Fortunately, such possibilities exist, e. g. , a hidden QCD-like sector based on the gauge group G2 Holland:2003jy ; Maas:2012wr ; Wellegehausen:2013cya . It has already been shown that this theory can support pure compact stellar objects Hajizadeh:2017jsw , which could be interpreted as pure dark stars. Already the one-flavor theory provides a conserved quantum number carried by the fermionic dark baryons Maas:2012wr , and thereby offers a dark matter candidate, alongside a rich spectrum of further states Wellegehausen:2013cya . The two-flavor case furthermore offers dark pions as additional stable states with a Wess-Zumino-Witten-like $3\to 2$ process, relevant to the dark matter relic density generation mechanism Hochberg:2014dra ; Hochberg:2014kqa ; Kulkarni:2022bvh . There is also an additional $3\to 2$ process which converts dark pions into dark fermionic baryons, and allows to build Fermi pressure, from accrued bosonic dark matter, required for neutron star stabilization Hajizadeh:2017jsw . G2-QCD therefore provides an ideal testbed for the previously mentioned challenge.
@@
For ordinary matter, we use a selection of the currently favored equations of state Kurkela:2014vha ; Dengler:2021qcq ; Pitz:2024xvh covering a wide range of phenomenology, and as much as possible data-driven constraints. This allows us to minimize the model-dependency of our investigation. The results will therefore only be quantitatively relevant if either we keep the amount of dark matter limited such that, within errors, neutron star properties remain the same. Or we stipulate that most neutron stars do not contain large amounts of dark matter, but some do, giving a distinct signal differing from ordinary neutron stars when observed. Alternatively, one can interpret our injection of G2-QCD equation of state seriously and demand modification of the ordinary matter equation of state to compensate for large modifications in the neutron star properties. However, this is beyond the scope of the present exploratory work, especially in view of the restricted set of available equations of state for the G2-QCD case. We will also assume throughout that the interactions between dark matter and ordinary matter, even at large densities of both, are negligible. This may be justified given that in most strongly-interacting dark matter scenarios the portal couplings to the SM are highly suppressed Hochberg:2015vrg ; Berlin:2018tvf ; Pomper:2024otb ; Bernreuther:2023kcg .
For the sake of completeness, we discuss briefly our choices of the equations of state in section 2, together with a brief overview of the salient features of G2-QCD. Our approach to building a two-component neutron star is standard Hajkarim:2024ecp , and briefly rehearsed in section 3. In section 4 we provide results for a range of dark matter masses and dark matter content. We provide the mass-radius relations, together with maximal and minimal values for both. We also provide the tidal deformability, which is the primary input to the gravitational wave signal for the merging of two neutron stars. Finally, taking our inputs at face value, we interpret our result in terms of limits on dark matter properties or dark matter contents in neutron stars. This part should be considered rather a blueprint than a hard statement, given our limited inputs, but can be improved arbitrarily by improving these inputs. Additional results are available in appendices 6.2 and 6.3.
@@ 2 Equations of state
Figure 1: The equations of state from Kurkela:2014vha and Wellegehausen:2013cya used in this work. The dark matter candidate mass is chosen to be the mass of the neutron for this plot for comparison. The end of the lines indicate the maximum value for the pressure for the respective equation of state used in this work. The black line indicates $\varepsilon=p$. For the ordinary matter equation of state, the stars denote the results from NChPT while the dots show the endpoints of the piecewise polytropes.
The equation of state is the key ingredient for the Tolman-Oppenheimer-Volkoff (TOV) equation Oppenheimer:1939ne which are commonly used to calculate macroscopic neutron star observables. To describe a neutron star consisting of two fluids we need the equations of state of both of them. A novelty of this work is the use of an equation of state of a confining, asymptotically free theory obtained from first principles. This is crucial as strongly-interacting dark matter models make use of gauge theories often without investigating its non-perturbative signatures. We want to complement this by the use of equations of states for ordinary matter in as much a model agnostic way that is able to meet the observational constraints. In this way we minimize the number of free parameters. We emphasize that this choice is not sufficient to cover all possible scenarios for neutron stars, as equations of states with (first order) phase transitions or twin-stars, for example, are not covered Mendes:2024hbn ; Christian:2023hez .
For both ordinary and dark matter, we will use polytropes to interpolate between data points. In Kurkela:2009gj , this was done to interpolate between different physical regimes. We use the same method to interpolate between data points, also for both the interpolation of the NChPT data points and the lattice data points. Details of this procedure are given in appendix 6.1.
@@ 2.1 Ordinary matter
For ordinary matter we work with the equations of state from Kurkela:2014vha . This work uses piecewise polytropes, to interpolate between a low density regime described by nuclear chiral perturbation theory (NChPT) Hebeler_2013 and a high density regime described by perturbative QCD Kurkela:2009gj . Equations of state that do not meet the observational constraints from NICER data on the mass, radius and the tidal deformability are omitted.
For this study, we use 3 limiting equations (EoS I, II and III) of state from Kurkela:2014vha . EoS I is the softest of the three, barely reaching the two solar mass limits. EoS II yields a maximal maximum mass. EoS III is the stiffest and while it might be too stiff to describe an ordinary neutron star, we still include it in our analysis as the addition of dark matter to a neutron star can result in smaller masses and radii. Therefore it might be able to meet observational constraints when considering mixed neutron stars. For the same reason we will employ central pressures larger than the ones stated in Kurkela:2014vha that result in the maximal mass. We will see that the addition of dark matter allows for stable solutions at larger central pressures Kain:2021hpk ; Valdez-Alvarado:2012rct . Figure 1 shows the different equations of state alongside the equation of state used to describe dark matter.
We extend the equations of state below the NChPT datapoints in Kurkela:2014vha by interpolating from zero using another polytrope with adiabatic index $\Gamma=5/3$. This corresponds to a free fermi gas and coincides with the equation of state $n\propto(\mu^{2}-m_{n}^{2})^{3/2}$ for low densities. We do not include a specified equation of state for the crust which is sufficient for the scope of this work, as the crust would hardly affect the radius and the effect on the mass would be negligible.
@@ 2.2 G2-QCD
Details on the equation of state, its determination and properties, can be found in  Wellegehausen:2013cya ; Maas:2012wr ; Hajizadeh:2017jsw . Here, we concentrate on the equation of state at zero temperature, which has been discussed in Wellegehausen:2013cya ; Hajizadeh:2017jsw . Of course, the lattice data is only available at discrete values of the dark baryonic chemical potential. The interpolation between data points is, as for the ordinary matter equation of state, performed using piecewise polytropes to ensure thermodynamic consistency.
Just like for the ordinary matter equation of state we use choose $\Gamma=5/3$ for the polytrope describing the low density behavior. This coincides with $n=c(\mu^{2}-m_{C}^{2})^{3/2}$ for low densities, where $m_{C}$ is the mass of the dark matter candidate and c is a free parameter as was considered in Hajizadeh:2017jsw . For the light ensemble and at high densities, we extrapolate the data using the fit form $n(\mu)=n_{s}/(\exp(a-b\mu)+1)$ fitted to the last few data points Wellegehausen:2013cya ; Hajizadeh:2017jsw . Herein $n_{s}$ is the lattice saturation density, i. e. the situation where Pauli blocking obstructs larger densities. This is a pure lattice artifact but the densities required in the following, as has already been observed previously Hajizadeh:2017jsw , are small enough for this to be not a relevant issue. In the heavy ensemble, the extrapolation region is never probed. The interpolation and extrapolation captures the main physical features of the lattice results while providing access to a larger range in the density. Note that, in contrast to the low-density case, the high-density fit function has not been chosen on physical grounds, but for its ability to describe the lattice data very accurately.
The equation of state has been obtained for two different values of the bare lattice parameters [^2] The only free parameter in the theory for the dark matter is the bare fermion mass., which will result in different mass spectra. The lightest particle in the theory is the pseudo-scalar dark pion which is unstable to potential decays to the standard model via messengers. The dark matter candidate is the lightest fermionic particle. In the two resulting ensembles the mass of the dark pion is 0.25 $m_{c}$ and 0.33 $m_{c}$ respectively. The original choice of the lattice simulation parameters were for a different reasons than dark matter Maas:2012wr , and thus do not apply here. It would require a major effort and incur computational costs of the order of a few tens of millions of core hours to provide further values  Wellegehausen:2013cya . This would only be justified by concrete observational hints for a particular mass value. However, for our exploratory investigations, the precise value does not matter too much. As will be seen, the impact is only a limited quantitative one, and the change between the ordinary matter EoSs has already a more substantial effect.
The resulting dark matter equation of states are also shown in figure 1.
@@ 3.1 Two-fluid TOV
The coupled two-fluid TOV equations can be solved using the equations of state of the two fluids. Iterating over the potentially nonidentical central pressures, one obtains the neutron star properties for each combination of central pressures ($p_{0,O}$ and $p_{0,D}$) by integrating the equations until the pressure vanishes. The radius at which the ordinary/dark matter pressure drops to zero within numerical precision is labeled ROR{O}/RDR{D}. $M_{O}$$M_{D}$ are the integrated masses of the two fluids respectively and $M_{tot}=M_{O}+M_{D}$. The solutions can result in a neutron star with a dark core ($r_{O}>r_{D}$) or a dark halo ($r_{O}<r_{D}$), which potentially also affects the gravitational wave signal of an inspiral of two neutron stars Ivanytskyi:2019wxd .
Dimensional analysis shows that the mass and the radius scale like $1/m_{N}^{2}$. Given an equation of state that can be trivially scaled by the mass of its constituents $m_{const}$, as is the case with our dark matter equation of state, halving the mass of the constituents would result in a star with four times the mass and four times the radius. To some extent, this scaling argument is still present in the two fluid case. A lighter dark matter candidate mass will result in a stronger impact of dark matter on the star. In the limiting cases where the central pressure of one of either the dark or ordinary matter is negligible ($p_{0,D}\gg p_{0,O}$ or $p_{0,D}\ll p_{0,O}$) this scaling argument is perfectly recovered as very heavy dark matter candidates will have no impact on the observables while a light dark matter candidate will result in a compact object dominated by dark matter that can eventually not be interpreted as a neutron star anymore.